Sol Summary – May 23rd

Crew 281 Sol Summary Report 23 May 2023
Sol: 9
Summary Title: Moon and Mars
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
Today an all-female EVA crew from Team Pegasus successfully completed a long EVA to explore the Overlook region. There was a lot of planning and contingencies that the crew discussed, especially since there were rains and storms last night. The crew were well prepared and narrowed down Mission Support efforts to specific location/coordinates per time estimations for each spot. They started out at 0758 for the intersection of Cow Dung and Hwy 1572, which was previously termed as White Moon. The area was interesting due to the geological features that made a unique mix as if they were from moon and mars. The crew witnessed shades and combinations of red and white at various locations around, a beautiful sight especially with the sunlight going in and out due to overcast weather. On arrival, Ana (Crew Scientist) collected rock and soil samples and Ritu (XO, EVA Lead) deployed drone for aerial footage. For the next stop, the crew wrapped up and drove on Hwy 1572 until they arrived at the ‘Y’ fork at the Overlook region. Weather was perfect, bright and sunny, there were more samples and some great drone shots collected.
The crew then left for Pooh’s Corner on their way back and acquired vertical and oblique drone images and videos for previously done geotechnical survey. This will aid to georeferencing the region under study. The crew left for MDRS right after and arrived at 1037.
During the EVA Comms did not work once the EVA crew went past Pooh’s corner. At the Hab Rachel and KC managed to set up the SDR to pick up the EVA Crew’s transmissions.
After the EVA, the crew grabbed a snack and jumped in on report writing and completed mid-mission video. KC (Crew Engineer) made a hearty veggie soup with handmade noodles. Which gave the crew an impetus to jump in the next big activity for the day, making STEM educational videos for waves and propagation in the Science dome. Megan (Commander) and Rachel (HSO) directed and shot the movies respectively, and the rest of the crew helped with acting and set up.
Green hab got very hot up to 43 C today, Megan watered and checked on the plants. KC did the routine hab operational checks. Rachel made more contacts on the ham radio station.
Dinner was finishing up the leftovers.
Look Ahead Plan: Tomorrow we will do 2-person EVA near the hab to check the antenna and conduct a drone payload delivery test. The drone will be launched from the observatory and flown to the crew on EVA. First test will be without a payload second with a payload.
There were no operations in the RAM today.
Anomalies in work: None
Weather: Winds and thunder
Crew Physical Status: Nominal
EVA: Completed 1 today

Reports to be filed:
Greenhab Report
EVA Report
EVA Request
Ops Report
Journalist Report
Photo of the Day

Support Requested:
Mission support is requested in case of storms if the antenna needs to come down or be retrieved.

Copyright © The Mars Society. All rights reserved. | Main Site